# Need for zkSpeed: Accelerating HyperPlonk for Zero-Knowledge Proofs

### v1

@@ 6.2. Benchmarks
HyperPlonk was evaluated using mock circuit workloads (Chen et al., 2022), as there is no publicly available compiler to generate real workloads. We similarly use synthetically generated workloads to model the performance of our architecture, which is standard for ZKP benchmarks, as performance primarily depends on the size of the workload. Similarly, GZKP (Ma et al., 2023) uses synthetic workloads to benchmark their implementation for workload sizes of and higher. NoCap (Samardzic et al., 2024) uses workloads from libsnark (lib, 2018) that prove relatively small circuits and scales them up to larger problem sizes because their proving times are dominated by fixed overheads on smaller ZKP circuits. In the context of HyperPlonk, workload statistics primarily affect the witness commit (sparse MSM) step. The scalar distributions of the MSMs in the Wiring Identity and Polynomial Opening are random because they are constructed in part from SHA3 challenges. Therefore, while the MSMs are data-dependent, they are not workload dependent and their runtimes are roughly the same at iso-problem size. All other steps in HyperPlonk are data-oblivious. From prior work (Daftardar et al., 2024; Ma et al., 2023; Zhang et al., 2021; lib, 2018), we know that in sparse MSMs, the scalars are typically 5-10% dense (i.e., full bit-width), and 95-99% sparse. Since the dense components of sparse MSMs are still runtime dominant, we assume a pessimistic upper bound of 10% dense scalars and 90% sparse scalars, of which 45% are 1s and 45% are 0s. The rest of the protocol steps operate on 100% dense scalars/MLE values. In our evaluation, we use five workloads from prior work (Chen et al., 2022; lib, 2018), and show our results in Table 3.
@@ 7.1. Pareto Space Analysis
Figure 9 shows the design space for a problem size of gates under four bandwidth scenarios. We sweep all the parameters in Table 2 and obtain Pareto curves for each bandwidth individually, and then construct the global Pareto curve from these four local Pareto curves. The key highlight from this plot is that HBM3-scale bandwidths (e.g. 1-4 TB/s (hbm, 2023)) do yield significant performance gains over, e.g., an HBM2-scale bandwidth (0.5 TB/s (Inc., 2020)). Beyond 300 mm2, the globally pareto-optimal design configurations yield over 2 speedups compared to 512 GB/s designs and over 700 speedups over the CPU baseline. This is because high-performance SumCheck designs quickly saturate 512 GB/s of bandwidth. In this analysis, we also include the cost of HBM PHYs (Samardzic et al., 2024; Ebel and Reagen, 2024; Mo et al., 2023; Inc., 2020; Samardzic et al., 2021, 2022; Kim et al., 2022a, b, 2023) where the PHY cost is 14.9 mm2 for a single HBM2 PHY and 29.6 mm2 for a single HBM3 PHY. For low-performance designs (below 100mm2), higher bandwidth becomes less effective because of the related PHY area costs. Figure 9 shows that zkSpeed remains viable even at bandwidths typical of DDR5 (Schlachter and Drake, 2019) (256 GB/s and below). While HBM allows exploring scalability benefits, less expensive memory technologies can be used to achieve Pareto-optimality within a target performance range (e.g., within 50 ms).
@@
| — | Bandwidth (GB/s) | 64, 128, 256, 512, 1T, 2T, 4T |
We analyze the area and runtime of selected points on the Pareto curve in Figure 10 to further understand bandwidth sensitivity. We pick Pareto points representing the highest-performing design point for each bandwidth level. In the area breakdown, moving from low to high-performance design points (A to D), the proportion of the SumCheck area increases significantly, because SumCheck is bandwidth-intensive, and higher bandwidth allows more parallel SumCheck PEs, boosting throughput. The MSM unit accounts for a large portion of the total area, but its absolute area remains unchanged when switching to high performance. This trend is also evident in the runtime breakdown: total runtime decreases as bandwidth increases, and the runtime contributions of the SumCheck-related processes (ZeroCheck, PermutationCheck, and OpenCheck) become smaller. Our analysis of the Pareto design points shows that high performance significantly depends on sufficient bandwidth, particularly improving the SumCheck computation. Conversely, for low-performance designs, the system utilizes less bandwidth and allocates more resources to MSM computation.
@@ 7.2. Bandwidth Sensitivity
Figure 12 shows how the speedups for MSM-related computation and SumCheck-related computation scale with increased PE count and bandwidth. These are two runtime-dominant components in our design points shown in Figure 10. We take the runtime of all MSM and SumCheck operations for 1 PE under 512 GB/s, and compute respective speedups to these numbers. Because MSMs are compute-bound, adding compute resources improves the speedups significantly, while adding bandwidth does not. We do not see perfectly linear speedup because of the serialization incurred in PolyOpen MSMs. SumChecks, which rely on a streaming-based approach, are memory-bound. As we add compute, we see linear speedups initially and then diminishing returns after saturating bandwidth. Consequently, in our pareto-optimal design space, we see most points along the curve outside high-performance regimes use at most 2 SumCheck PEs compared to 8 or 16 MSM PEs.
7.3. Iso-CPU-Area Comparisons
@@
After picking each design, we run synthetic benchmarks over problem sizes . Figure 11 shows the speedup of each design over CPU baseline, and the breakdown across different steps of the protocol to understand where our speedups come from. In general, we get more speedup from our MSM units than the SumCheck units. This intuitively follows, given our observations that MSMs are compute-bound and more robust to bandwidth constraints. Additionally, the CPU poorly handles sparse computations because it serially computes the point addition for -valued scalars. PolyOpen MSMs also incur serialization costs that we reduce by overlapping MSM executions where possible. The variations in speedups over different problem sizes are an artifact of our choice to highlight different Pareto points per problem size; for example, at size problems, a dual-core MSM is used, while at size a single-core MSM is chosen. This is because the area costs of on-chip MLE SRAM begins to dominate; choosing a weaker MSM in turn reduces achievable speedup. Storing MLE tables entirely off-chip may yield larger MSM speedups at higher bandwidth costs for SumChecks. These tradeoffs can be explored in future work.
| Workload | Problem Size | Runtime (ms) | |
| CPU | zkSpeed | ||
| Zcash | 1429 | 1.984 (720) | |
| Auction | 8619 | 11.405 (755) | |
| Rescue-Hash Invocations | 18637 | 18.335 (844) | |
| Zexe’s Recursive Circuit | 37469 | 43.451 (862) | |
| Rollup of 10 Pvt Tx | 74052 | 86.181 (859) | |
@@
run time breakdown.
utilization breakdown.
7.3.1. Runtime Breakdown and Utilization
Figure 13 shows the latency breakdown of HyperPlonk on a CPU and zkSpeed. The CPU executes kernels sequentially, enabling detailed profiling; we report step latency for zkSpeed due to its parallel scheduling of kernels. As expected, the majority of time goes to processing MSMs, while a handful of other kernels account for single percentage points of runtime. Figure 14 presents the utilization of each unit and relative (datapath) area allocation (design in Table 5). The utilizations vary from over 70% to 5% for some modules. zkSpeed was intentionally designed (via the design space search in Figure 9) to allocate resources to cores to optimize high performance per area, i.e. the Pareto front. This can be seen in our analysis in two ways. First, the cores taking up most area, notably MSM at 64.6%, are the most used, and following the profiling data (Figure 13) require the most speedup. Second, though some units are used infrequently, they (i) take up little area and (ii) are essential to accelerate to achieve the speedups desired (i.e., 2-3 orders of magnitude). For example, the SHA-3 unit is rarely used, but provides a speedup of over over the CPU and takes only area. Additionally, consider that MLE Combine makes up 3.3% of the CPU runtime, but without acceleration caps speedup to a mere 30.3, thus justifying its 5.85% area allocation and relatively low utilization.
@@ 7.4. Workload Evaluation
We pick a fixed design and show the end-to-end speedups in Table 3. As mentioned in Section 6.2, we assume a pessimistic 10% sparse scalar statistics for each workload. Our fixed design has one MSM unit with 9-bit windows, 16 PEs, and 2048 points per PE, with 1 FractionMLE PE, 2 SumCheck PEs, 11 MLE Update PEs, and 4 modular multipliers per MLE Update PE. The area is provided in Table 5. At roughly iso-CPU-core compute cost, zkSpeed achieves a geometric mean speedup of 801 over the CPU, with total area of 366.46 mm2, total average power of 170.88 W, and total power density of 0.46 W/mm2, which is within that of our CPU (Daftardar et al., 2024).
| Area (mm²) | Average Power (W) | |
@@ 8. Related Work
Much of the prior body of crytographic hardware and systems research has focused on Fully Homomorphic Encryption and Multi-Party Computation (Kim et al., 2022a, b, 2023; Samardzic et al., 2021, 2022; Soni et al., 2023; Mo et al., 2023; Garimella et al., 2023; Neda et al., 2024). ZKP hardware research is relatively newer, and has focused primarily on accelerating NTTs and MSMs (Liu et al., 2024a; Ji et al., 2024; Lu et al., 2022; Liu et al., 2024c; Zhou et al., 2024; Liu et al., 2024b; Qiu et al., 2024; Butt et al., 2024; Zhu et al., 2024; Chen et al., 2024; Wang and Gao, 2023; Yang et al., 2025; Wang and Gao, 2025; Hirner et al., 2025). A few recent works have accelerated SumChecks on GPU (Lu et al., 2024) and ASIC (Samardzic et al., 2024) as well as hashing alternatives to SHA-based hash functions (Sheybani et al., 2025; Ahmed et al., 2024; Wang and Gao, 2025). Some systems accelerate end-to-end Groth16 proofs (using NTTs and MSMs) on GPU (Ma et al., 2023) and ASICs (Daftardar et al., 2024; Zhang et al., 2021). SZKP is presently the only ASIC that accelerates Groth16 proofs entirely on-chip. NoCap (Samardzic et al., 2024) is an ASIC that accelerates the Spartan protocol, using Orion as the polynomial commitment scheme. NoCap focuses on SumCheck and NTTs used in Spartan. We compare zkSpeed with two ASICs that accelerate full proofs end-to-end.
SZKP is the state-of-the-art for accelerating Groth16 proofs, focusing on scalable MSM designs and (quasi)-deterministic scheduling for Pippenger’s algorithm. It accelerates all MSMs, including Sparse G2 MSMs, achieving geomean speedups of 493 over a CPU. SZKP improves on PipeZK (Zhang et al., 2021), the first hardware accelerator for Groth16 proofs. While Groth16 and HyperPlonk have similar application spaces, as mentioned in Section 1, the key advantage of using HyperPlonk is the universal setup, which means that the protocol parameters are application-agnostic. For Groth16, every new application that wants to use a ZKP needs its own trusted setup ceremony (cer, 2023), which is impractical as the application space grows. Given this context and the recent shift away from Groth16 (Nelson, 2022), the slightly larger proof sizes are considered a reasonable tradeoff.
@@
For easier comparison, Table 4 compares zkSpeed, NoCap, and SZKP’s protocols and software and hardware costs. zkSpeed’s parent Hyperplonk has the slowest software prover, reflecting the complexity of the protocol. Of note, Spartan’s prover is slow; NoCap’s authors explain this is due to inefficient implementation.
We compare NoCap’s hardware implementation using the design point and numbers from their paper scaled to 7nm using scale factors from prior work (Mo et al., 2023; Daftardar et al., 2024). We then select a zkSpeed configuration with roughly similar prover time. At iso-prover time, zkSpeed incurs a nearly area cost in return for a three orders-of-magnitude reduction in proof size. NoCap’s lower costs come from eliminating MSMs, having simpler sumchecks, and using a 64-bit Goldilocks-64 prime field that yields smaller modmuls. In contrast, zkSpeed supports arbitrary 255-bit and 381-bit primes for MLEs and elliptic curves points, respectively. Consequently, NoCap runs all operations several times, including SumChecks 3 times, to obtain 128 bits of security. We further compare zkSpeed with an iso-area SZKP (Groth16) implementation, giving them the benefit of zkSpeed’s improved MSMs, and optimistically scale up their design to use the BLS12-381 curve. This design, SZKP+, enjoys a 6 reduction in proving time compared to zkSpeed, largely because it has fewer MSMs on its critical path. These speedups come at the cost of circuit-specific setup, incurring large costs any time the application is updated. In sum, NoCap, SZKP, and zkSpeed each address different application domains, representing a range of trade-offs ranging from security and protocol properties to software/hardware costs.
@@ 9. Conclusion
This paper presents zkSpeed, the first work to accelerate HyperPlonk proofs in hardware, which offers time complexity compared to prominent zkSNARKs that rely on computational primitives that have complexity (e.g. Groth16). zkSpeed constitutes accelerator units for all core HyperPlonk functions, with special attention paid to prominent kernels: SumCheck and MSM. zkSpeed is a modular architecture, and we leverage a performance model to conduct design space optimization, and analyze the pareto frontier to identify well performing designs. A zkSpeed accelerator with 366 mm2 and 2 TB/s of bandwidth achieves gmean speedup of over CPU baselines, demonstrating the promise zkSpeed offers to accelerate HyperPlonk.

### v2

@@ 6.2. Benchmarks
HyperPlonk was evaluated using mock circuit workloads (Chen et al., 2022), as there is no publicly available compiler to generate real workloads. We similarly use synthetically generated workloads to model the performance of our architecture, which is standard for ZKP benchmarks, as performance primarily depends on the size of the workload. Similarly, GZKP (Ma et al., 2023) uses synthetic workloads to benchmark their implementation for workload sizes of and higher. NoCap (Samardzic et al., 2024) uses workloads from libsnark (lib, 2018) that prove relatively small circuits and scales them up to larger problem sizes because their proving times are dominated by fixed overheads on smaller ZKP circuits. In the context of HyperPlonk, workload statistics primarily affect the witness commit (Sparse MSM) step. The scalar distributions of the MSMs in the Wiring Identity and Polynomial Opening are random because they are constructed in part from SHA3 challenges. Therefore, while all MSMs are data-dependent, the overall runtimes are effectively workload-agnostic and roughly the same at iso-problem size. All other steps in HyperPlonk are data-oblivious. From prior work (Daftardar et al., 2024; Ma et al., 2023; Zhang et al., 2021; lib, 2018), we know that in Sparse MSMs, the scalars are typically 5-10% dense (i.e., full bit-width), and 95-99% sparse. Since the dense components of Sparse MSMs are still runtime dominant, we assume a pessimistic upper bound of 10% dense scalars and 90% sparse scalars, of which 45% are 1s and 45% are 0s. The rest of the protocol steps operate on 100% dense scalars/MLE values. In our evaluation, we use five workloads from prior work (Chen et al., 2022; lib, 2018), and show our results in Table 3.
@@ 7.1. Pareto Space Analysis
Figure 9 shows the design space for a problem size of gates under seven bandwidth scenarios. We sweep all the parameters in Table 2 and obtain Pareto curves for each bandwidth individually, and then construct the global Pareto curve from these seven local Pareto curves. The key highlight from this plot is that HBM3-scale bandwidths (e.g., 1-4 TB/s (hbm, 2023)) do yield significant performance gains over, e.g., an HBM2-scale bandwidth (0.5 TB/s (Inc., 2020)). Beyond 300 mm2, the globally Pareto-optimal design configurations yield over 2 speedups compared to 512 GB/s designs and over 700 speedups over the CPU baseline. This is because high-performance SumCheck designs quickly saturate 512 GB/s of bandwidth. In this analysis, we also include the cost of HBM PHYs (Samardzic et al., 2024; Ebel and Reagen, 2024; Mo et al., 2023; Inc., 2020; Samardzic et al., 2021, 2022; Kim et al., 2022a, b, 2023) where the PHY cost is 14.9 mm2 for a single HBM2 PHY and 29.6 mm2 for a single HBM3 PHY. For low-performance designs (below 100mm2), high-bandwidth memory becomes less effective, as these designs tend to underutilize the available bandwidth while still incurring the high PHY area overhead associated with HBM-scale memory. Figure 9 shows that zkSpeed remains viable even at bandwidths typical of DDR5 (Schlachter and Drake, 2019) (256 GB/s and below). While HBM allows exploring scalability benefits, less expensive memory technologies can be used to achieve Pareto-optimality within a target performance range (e.g., within 50 ms).
@@
| Memory | Bandwidth (GB/s) | 64, 128, 256, 512, 1T, 2T, 4T |
We analyze the area and runtime of selected points on the Pareto curve in Figure 10 to further understand bandwidth sensitivity. We pick Pareto points representing the highest-performing design point for each bandwidth level. In the area breakdown, moving from low to high-performance design points (A to D), the proportion of the SumCheck area increases significantly, because SumCheck is bandwidth-intensive, and higher bandwidth allows more parallel SumCheck PEs, boosting throughput. The MSM unit accounts for a large portion of the total area, but its absolute area remains unchanged when switching to high performance. This trend is also evident in the runtime breakdown: total runtime decreases as bandwidth increases, and the runtime contributions of the SumCheck-related processes (ZeroCheck, PermCheck, and OpenCheck) become smaller. Our analysis of the Pareto design points shows that high performance significantly depends on sufficient bandwidth, particularly improving the SumCheck computation. Conversely, for low-performance designs, the system utilizes less bandwidth and allocates more resources to MSM computation.
@@ 7.2. Bandwidth Sensitivity
Figure 11 shows how the speedups for MSM-related computation and SumCheck-related computation scale with increased PE count and bandwidth. These are two runtime-dominant components in our design points shown in Figure 10. We take the runtime of all MSM and SumCheck operations for 1 PE under 512 GB/s, and compute respective speedups to these numbers. Because MSMs are compute-bound, adding compute resources improves the speedups significantly, while adding bandwidth does not. We do not see perfectly linear speedup because of the serialization incurred in Polynomial Opening MSMs. SumChecks, which rely on a streaming-based approach, are memory-bound. As we add compute, we see linear speedups initially and then diminishing returns after saturating bandwidth. Consequently, in our Pareto-optimal design space, we see most points along the curve outside high-performance regimes use at most 2 SumCheck PEs compared to 8 or 16 MSM PEs.
7.3. CPU Performance Comparisons
@@
7.3.1. Runtime Breakdown and Utilization
Figure 12 shows the latency breakdown of HyperPlonk proving on a CPU and a pareto-optimal zkSpeed design for a problem size of gates. The CPU executes kernels sequentially, enabling detailed profiling; we report step latency for zkSpeed due to its parallel scheduling of kernels. As expected, the majority of time goes to processing MSMs, while a handful of other kernels account for single percentage points of runtime. Figure 13 presents the utilization of each unit and relative (datapath) area allocation (design in Table 5). The utilizations vary from over 70% to 5% for some modules. zkSpeed was intentionally designed (via the design space search in Figure 9) to allocate resources to cores to optimize high performance per area, i.e. the Pareto front. This can be seen in our analysis in two ways. First, the cores taking up most area, notably MSM at 64.6%, are the most used, and following the profiling data (Figure 12) require the most speedup. Second, though some units are used infrequently, they (i) take up little area and (ii) are essential to accelerate to achieve the speedups desired (i.e., 2-3 orders of magnitude). For example, the SHA-3 unit is rarely used, but provides a speedup of over over the CPU and takes only area. Additionally, consider that MLE Combine makes up 3.3% of the CPU runtime, but without acceleration caps speedup to a mere 30.3, thus justifying its 5.85% area allocation and relatively low utilization.
run time breakdown.
utilization breakdown.
7.3.2. Iso-CPU Area Comparisons
After picking each Pareto-optimal design, we run synthetic benchmarks over problem sizes . Figure 14 shows the speedup of each design over CPU baseline, and the breakdown across different steps of the protocol to understand where our speedups come from. In general, we get more speedup from our MSM units than the SumCheck units. This intuitively follows, given our observations that MSMs are compute-bound and more robust to bandwidth constraints. Additionally, the CPU poorly handles sparse computations because it serially computes the point addition for -valued scalars. Polynomial Opening MSMs also incur serialization costs that we reduce by overlapping MSM executions where possible. The variations in speedups over different problem sizes are an artifact of our choice to highlight different Pareto points per problem size; for example, at size problems, a dual-core MSM is used, while at size a single-core MSM is chosen. This is because on-chip MLE SRAM area begins to dominate, limiting MSM compute area and therefore achievable speedup. Storing MLE tables entirely off-chip may improve MSM speedups at higher SumCheck bandwidth costs. These tradeoffs can be explored in future work.
| Workload |
Runtime (ms) | |||
| CPU | zkSpeed | ||||
| Zcash | 1429 | 1.984 (720) | |||
| Auction | 8619 | 11.405 (755) | |||
| Rescue-Hash Invocations | 18637 | 22.082 (844) | |||
| Zexe’s Recursive Circuit | 37469 | 43.451 (862) | |||
| Rollup of 10 Pvt Tx | 74052 | 86.181 (859) | |||
@@ 7.4. Workload Evaluation
We pick a fixed design and show the end-to-end speedups in Table 3. As mentioned in Section 6.2, we assume pessimistic 10% sparse scalar statistics for each workload. Our design has one MSM unit with 9-bit windows, 16 PEs, and 2048 points per PE, with 1 FracMLE PE, 2 SumCheck PEs, 11 MLE Update PEs, and 4 modular multipliers per MLE Update PE. The area breakdown is provided in Table 5. At roughly iso-CPU-core compute cost, zkSpeed achieves a geometric mean speedup of 801 over the CPU, with total area of 366.46 mm2, total average power of 170.88 W, and total power density of 0.46 W/mm2, which is within that of our CPU (Daftardar et al., 2024).
| Module | Area (mm²) | Average Power (W) |
@@ 8. Related Work
Much of the prior body of cryptographic hardware and systems research has focused on Fully Homomorphic Encryption and Multi-Party Computation (Kim et al., 2022a, b, 2023; Samardzic et al., 2021, 2022; Soni et al., 2023; Mo et al., 2023; Garimella et al., 2023; Neda et al., 2024; Ebel et al., 2025; Jha and Reagen, 2024). ZKP hardware research is relatively newer, and has focused primarily on accelerating NTTs and MSMs (Liu et al., 2024a; Ji et al., 2024; Lu et al., 2022; Liu et al., 2024c; Zhou et al., 2024; Liu et al., 2024b; Qiu et al., 2024; Butt et al., 2024; Zhu et al., 2024; Chen et al., 2024; Wang and Gao, 2023; Yang et al., 2025; Wang and Gao, 2025; Hirner et al., 2025; Zhang and Franchetti, 2025). A few works have accelerated SumChecks on GPU (Lu et al., 2024) and ASIC (Samardzic et al., 2024) as well as hashing alternatives to SHA-based hash functions (Sheybani et al., 2025; Ahmed et al., 2024; Wang and Gao, 2025). Some systems accelerate end-to-end Groth16 proofs (using NTTs and MSMs) on GPU (Ma et al., 2023) and ASICs (Daftardar et al., 2024; Zhang et al., 2021), while others accelerate Spartan proofs with Orion commitments (using NTTs and SumChecks) (Samardzic et al., 2024). We compare zkSpeed with two ASICs, SZKP and NoCap, that accelerate full proofs end-to-end.
SZKP is the state-of-the-art for accelerating Groth16 proofs, focusing on scalable MSM designs and (quasi)-deterministic scheduling for Pippenger’s algorithm. It accelerates all MSMs, including Sparse G2 MSMs (not present in HyperPlonk), achieving geomean speedups of 493 over a CPU. SZKP improves on PipeZK (Zhang et al., 2021), the first hardware accelerator for Groth16 proofs. While Groth16 and HyperPlonk have similar application spaces, as mentioned in Section 1, the key advantage of using HyperPlonk is the universal setup, which means that the protocol parameters are application-agnostic. For Groth16, every new application that wants to use a ZKP needs its own trusted setup ceremony (cer, 2023), which is impractical as the application space grows. Given this context and the recent shift away from Groth16 (Nelson, 2022), the slightly larger proof sizes are considered a reasonable tradeoff.
@@
Comparison: Table 4 compares zkSpeed, NoCap, and SZKP’s protocols, software, and hardware costs. zkSpeed’s parent HyperPlonk has the slowest software prover, reflecting the complexity of the protocol. Of note, Spartan’s prover is slow; NoCap’s authors explain this is due to inefficient implementation.
We compare NoCap’s hardware implementation using the design point and numbers from their paper scaled to 7nm using scale factors from prior work (Mo et al., 2023; Daftardar et al., 2024). We then select a zkSpeed configuration with roughly similar prover time to NoCap. At iso-prover time, zkSpeed incurs a nearly area cost in return for a three orders-of-magnitude reduction in proof size. NoCap’s lower costs come from not having MSMs, having simpler SumChecks, and using a 64-bit Goldilocks-64 prime field that yields smaller modmuls. Consequently, NoCap runs all operations several times, including SumChecks 3 times, to obtain 128 bits of security. In contrast, zkSpeed supports arbitrary 255-bit and 381-bit primes for MLEs and elliptic curves points, respectively. We further compare zkSpeed with an iso-area SZKP (Groth16) implementation, giving them the benefit of zkSpeed’s improved MSMs, and optimistically scale up their design to use the BLS12-381 curve. This design, SZKP+, enjoys a 6 reduction in proving time compared to zkSpeed, largely because it has fewer MSMs on its critical path. These speedups come at the cost of circuit-specific setup, incurring large costs any time the application is updated. In sum, NoCap, SZKP, and zkSpeed address different application domains, representing a range of tradeoffs ranging from security properties to software/hardware costs.
@@ 9. Conclusion
This paper presents zkSpeed, the first work to accelerate HyperPlonk proofs in hardware, which offers time complexity compared to prominent zkSNARKs that rely on computational primitives that have complexity (e.g. Groth16). zkSpeed constitutes accelerator units for all core HyperPlonk functions, with special attention paid to prominent kernels: SumCheck and MSM. zkSpeed is a modular architecture, and we leverage a performance model to conduct design space optimization, and analyze the pareto frontier to identify well performing designs. A zkSpeed accelerator with 366 mm2 and 2 TB/s of bandwidth achieves geomean speedup of over CPU baselines, demonstrating the promise zkSpeed offers to accelerate HyperPlonk.